d-1. Determine the cost per unit output for part b. 1.54  

d-2. Determine the cost per unit output for part c. 1.50

e. The product is sold at $6 per unit. Assume that the cost of a drilling machine (fixed cost) is $34,000 and the company produces 7,400 units per week. Assume that four drilling machines are used for production. If the company had an option to buy the same part at $5 per unit, what would be the break-even number of units? (In your calculations, use the two-digit cost per unit from page d-1. Round your answer to the nearest whole number.)

E.___________

d-1 and d-2 are correct, it says you need it so I included it.
Answer #1

As per Chegg Guidelines we cannot provide direct answers. However, I've answered it as below
Violating them will get my account suspended.
So please understand and like this answer

Fixed Cost = No. of drilling machines used x Cost per machine = 4 x 34,000 = 136,000 = Total Fixed Cost

Total Variable Cost = Price x No. Of units = 5 x 7,400 = 37,000

Total Revenue = 6 x 7,400 = 44,400

BEP = Fixed Cost / (Revenue - Variable Cost) = 136,000 / (44,400 - 37,000) = 136,000 / 7,400 = 18.37837837837838‬

or 18 units